11. Wodera cooperative society of farmers has 50 hectare of land to grow two crops X and Y. The profit from crops X and Y per hectare are estimated as Birr 10,500 and Birr 9,000 respectively. To control weeds, a liquid herbicide has to be used for crops X and Y at rates of 20 liters and 10 liters per hectare. Further, no more than 800 liters of herbicide should be used in order to protect fish and wild life using a pond which collects drainage from this land. How much land should be allocated to each crop so as to maximize the total profit of the society? (Solve By graphic method).

1
Expert's answer
2021-12-31T13:05:38-0500

profit:

"Z=10500x_1+9000x_2"

total number of hectares:

"x_1+x_2\le50"

amount of of herbicide:

"20x_1+10x_1\le 800"

x1 is number of hectares for crop X,

x2 is number of hectares for crop Y




for Extreme Points:

"x_1+x_2=50"

"x_1=0\implies x_2=50"


"20x_1+10x_2=800"

"x_2=0\implies x_1=40"


for intersection of graphs:

"x_1=50-x_2"

"2(50-x_2)+x_2=80"

"x_2=20,x_1=30"


Objective function values at extreme points:

"Z(40,0)=420000"

"Z(0,50)=450000"

"Z(30,20)=495000"


The maximum value of the objective function Z=495000 occurs at the extreme point (30,20).

Hence, the optimal solution to the given LP problem is : x1=30, x2=20 and max Z=495000